Use Slater's rules to estimate Zeff for the titanium atom and vanadium atom.
Answer #1

IS2 262 2p 3s 3p 3d 452 Ti 6=(Ix0.35)+ (tox o. 85) + Croxi) 0.35+ 2.5+ 1D Cn-25 (n-i9 S8.8 3. 15 8.85+ 10 = 22-18-2S = Z= ffthank you vmuch...